# OmniGen: Unified Image Generation

### v1

@@ 3.2.3 Computer Vision Tasks
We introduce classic computer vision tasks to enhance the image generation capabilities of the model. For low-level vision tasks (low-light image enhancement [66], deraining [71], deblurring [46], inpainting [53], outpainting [53] and colorization [58]), where the annotation itself is an image, we only add text instructions, which were randomly sampled from instructions generated by GPT-4o. For high-level tasks, we choose to represent all annotations as images. We used LAION [58] as the source image and annotations from [53] as the target to construct image pairs (such as source image and its human pose mapping). The annotations include human pose, depth mapping, canny, and segmentation. Additionally, we also use several datasets for referring image segmentation, including RefCOCO [27], ADE20k [76], and ReasonSeg [30]. As shown in Figure 3-(c), the input is the source image and a natural language expression, the output is an image with the corresponding object highlighted in blue.
The purpose of constructing these datasets is not merely to endow the model with these functionalities. We also aim to transfer the knowledge acquired from these traditional computer vision tasks to image generation tasks, thereby achieving more sophisticated image generation capabilities. Our experiments have also demonstrated that multi-task learning enables the model to exhibit emergent abilities.
@@ 3.3 Few-shot to Image
We constructed a few-shot to image dataset to stimulate the model’s in-context learning capabilities. Specifically, for each task described in the preceding sections, we randomly selected a few examples and combined the original input with these examples to form new inputs. The specific data format can be referenced in Figure 3-(e). Due to limitations in training resources, we opted to use only one example to enhance training efficiency.
@@ 4 Experimental Results
In this section, we show the results of OmniGen in image generation tasks and traditional vision tasks.
@@ 4.1.1 Qualitative Results
Figure 6: The results of subject-driven generation. OmniGen can generate new image based on the objects from reference images. When the reference image contains multiple objects, OmniGen can automatically identify the required objects based on textual instructions.
Figure 7: The results of OmniGen in different image generation tasks.
Figure 5 shows the results of the text-to-image task. It can be observed that OmniGen effectively follows the textual descriptions to generate images with arbitrary aspect ratios.
Figure 6 presents the outcomes of the subject-driven generation task. Our model can extract the required objects from the given reference images and generate new images accordingly. Furthermore, when the reference image contains multiple objects, the model can directly select the needed objects based on textual instructions (e.g., the cat in the figure) without requiring additional preprocessing steps such as image cropping or face recognition.
Figure 7 summarizes the results of other image generation tasks, demonstrating that the model can handle various downstream tasks based on multi-modal instructions.
@@ 4.1.2 Text to Image
Following [13], we evaluate text-to-image generation capability of OmniGen on the GenEval [22] benchmark. We compared the performance of our model with the reported results of other popular image generation models, as summarized in Table 2. Surprisingly, our model achieved similar performance compared to the current state-of-the-art diffusion models, such as SD3, which underscores the effectiveness of our framework. The GenEval benchmark does not reflect the aesthetic quality of images, we will leave this aspect for future evaluation.
Notably, our model has only 3.8 billion parameters, whereas the SD3 model has a total of 12.7 billion parameters (more than three times that of ours). Current diffusion models typically adopt an encoder-decoder architecture, utilizing an additional encoder to encode textual conditions (this text encoder alone is larger than our entire model). In contrast, our model architecture is significantly simplified, eliminating the cost of an additional text encoder, thereby greatly enhancing the efficiency of parameter utilization. Besides, we employed only 0.1 billion image data, whereas SD3 used over 1 billion (more than ten times that of ours), highlighting the role of multitask data X2I in enhancing text-to-image capabilities.
@@
Model Params Data Overall object object Counting Colors Position binding
@@
IF-XL 5.5B+4.7B 1.2B 0.61 0.97 0.74 0.66 0.81 0.13 0.35
SD3 8.0B+4.7B 1.0B 0.68 0.98 0.84 0.66 0.74 0.40 0.43
OmniGen 3.8B 0.1B 0.70 0.99 0.86 0.62 0.85 0.32 0.54
Table 2: Results of GenEval Benchmark. means the parameter of the frozen text encoder. Our model achieves similar performance with a relatively small scale and limited data.
4.1.3 Image Edit
@@
InstructPix2Pix [4] 0.834 0.219 0.762
MagicBrush [72] 0.838 0.222 0.776
PnP [62] 0.521 0.089 0.153
Null-Text Inv. [44] 0.761 0.236 0.678
EMU-Edit [59] 0.859 0.231 0.819
OmniGen 0.836 0.233 0.804
Table 3: Results on EMU-Edit test data. As a universal model, OmniGen demonstrates performance comparable to that of the best proprietary models.
We compare OmniGen with other state-of-the-art image editing models on EMU-Edit [59] dataset, which includes seven different operations: background alteration, comprehensive image changes, style alteration, object removal, object addition, localized modifications, and color/texture alterations. We measure three metrics: 1) CLIP-I: CLIP image similarity between the source image and output image; 2) DINO: DINO similarity between the source image and output image; and 3) CLIP-T: CLIP text-image similarity between edited image and target caption. DINO and CLIP-I similarity scores measure the model’s ability to preserve elements from the source image, while CLIP-T measures how well the model followed the instructions. As shown in Table 2, our model significantly outperforms InstructPix2Pix [4], and exhibits comparable performance to the current state-of-the-art model: EMU-Edit [59].
4.1.4 DreamBooth
We evaluate the single-entity subject-driven generation capability on DreamBench [57]. The DreamBench contains 750 prompts for 30 subjects (e.g., dog and toy). For each prompt, we generate 4 images, resulting in a comprehensive evaluation set of 3,000 images. Following Kosmos-G [49], we only select one image as input from the 4-7 provided images for each subject. We adopted DINO and CLIP-I from DreamBooth to assess subject fidelity, and CLIP-T for text fidelity. All results are summarized in Table 4. Compared to methods based on fine-tuning, our approach maintains a comparable level of text fidelity while better preserving the subject from the source image. Compared with models without fine-tuning, OmniGen significantly outperforms both Re-Imagen and Kosmos-G, and demonstrates superior subject fidelity relative to SuTI.
@@ 4.1.4 DreamBooth
Fine-Tuning
Textual Inversion [15] 0.569 0.780 0.255
DreamBooth [57] 0.668 0.803 0.305
BLIP-Diffusion [32] 0.670 0.805 0.302
Test Time Tuning Free
Re-Imagen [8] 0.600 0.740 0.270
SuTI [10] 0.741 0.819 0.304
Kosmos-G [49] 0.694 0.847 0.287
OmniGen 0.801 0.847 0.301
Table 4: Results on DreamBench. OmniGen achieves better subject fidelity and text fidelity.
4.1.5 Visual Conditional Controls
Image-based prompts can provide detailed spatial conditioning controls for diffusion models. To evaluate this ability of OmniGen, we use the dataset and script from  [33]. This benchmark includes ADE20K test dataset for segmentation mask condition, and evaluation split of MultiGen-20M for canny edge map, hed edge map, and depth map condition. For each condition, the controllability is evaluated by measuring the similarity between the input conditions and the extracted conditions from generated images of diffusion models. The experimental results are shown in Table 5. We can find that our model achieves optimal results on segmentation mask and hed edge map conditions, and obtains competitive results for canny edge map and depth map conditions.
@@ 4.1.5 Visual Conditional Controls
Figure 8: The results of OmniGen in traditional various vision tasks.
Figure 9: Top: Comparison with ControlNet. ControlNet involves two steps: first, using the corresponding detector to extract information from the reference image, and then loading the appropriate ControlNet module to process the visual conditions. In contrast, OmniGen completes the entire task in a single step, achieving superior image quality. Bottom: Results using different visual conditions. It can be observed that more detailed conditions lead to a higher similarity between the model’s output and the reference image.
@@ 4.2 Computer Vision Tasks
We present several qualitative results of computer vision tasks in Figure 8. OmniGen can handle various low-level vision tasks such as deraining, deblurring, and inpainting. In the bottom of Figure 8, we can see that OmniGen is also able to handle high-level tasks, such as human pose recognition and depth estimation.
So far, we have demonstrated that our model can generate images well based on visual conditions while also extracting visual conditions from raw images. This motivates us to ponder: can we directly use the model to generate new images based on a reference image in only one step, instead of first using a processor to extract spatial condition information and then inputting it into the model for generation? Surprisingly, even without having encountered such a task before, OmniGen handles it admirably. As shown in Figure 9, the existing workflow for ControlNet involves using a detector to extract spatial condition information from the reference image, and then loading the corresponding control module to model the spatial condition information for image generation, which requires multiple network components and operations. Now, only based on our model, we can directly input the reference image and text instruction (e.g., Follow the depth mapping of this image <reference-image> to generate new image. The text description for new image is “…”) to generate an image in only one step without any additional intermediate procedures. It can be observed that the model comprehends the instruction well; when tasked with using the human pose from the reference image, it perfectly replicates the human pose, and when using depth mapping, it retains more details, such as the folds in clothing.
@@ 5 Further Analysis
LLMs demonstrate remarkable generalization capabilities, achieving impressive performance in previously unseen tasks and domains. Furthermore, they can boost performance through mechanisms such as in-context learning and chain of thought. We observe similar functionalities in OmniGen as well, and present our findings in this section.
@@ 5.1 Emgerging Capabilities
By standardizing all tasks into a unified format and training on X2I dataset, OmniGen can acquire universal knowledge and allow knowledge transfer across different scenarios and tasks, thus enabling the generation capabilities on unseen tasks and domains. We illustrate several emerging capabilities using the following examples.
@@
Task Composition. In real-world applications, user requirements often involve combinations of tasks. As shown in Figure 10-(a), our model is capable of simultaneously processing multiple instructions, including those for different tasks (Image inpainting and change the color of hair to white) as well as multiple instructions for the same task (Add a sunglasses to the man’s face, and change the color of clothes to blue). These results highlight our model’s versatility and potential for widespread adoption in the wild.
Implicit Combination of Tasks. In addition to explicit task combinations, our model is capable of performing multiple tasks implicitly through a single instruction. As demonstrated in Figure 9, upon receiving the input like “follow the human pose/depth mapping to generate the image: …”, our model can extract the relevant conditional information (such as human pose, depth mapping, etc.) from the reference image and generate a new image based on the captured condition. This process is implicit, with all processing completed internally within the model, thus only requiring the user to input a simple command. This negates the need for explicit conditional extraction using other models prior to input into the diffusion process, as is necessary with existing systems like ControlNet [73].
In-context Learning for Unseen Tasks and Domains. As illustrated in Figure 10-(b), by providing an example, the model is capable of successfully completing a novel task: generating images based on provided scribble data, which is not encountered during training. To explore whether in-context learning can boost existing abilities on new domains, we show several examples from the FSS [34] dataset, which contains objects that have never been seen or annotated in previous datasets. In the left of Figure 10-(c), we can see that OmniGen is not familiar with the concepts of “pencil sharpeners” and “chess queens”, and it cannot identify them from images. However, when provided with an example, the model is capable of making accurate predictions, demonstrating that in-context learning can enhance the model’s generalization ability across different domains.
End-to-end Workflow. Users typically need to load multiple models and perform multi-step processing to ultimately generate a satisfactory image, making the workflow very cumbersome and costly. This complexity has led to the development of open-source tools and pipelines like ComfyUI [^2] https://github.com/comfyanonymous/ComfyUI. Our model possesses both excellent multi-modal understanding and image generation capabilities, enabling it to complete a lot of tasks without relying on external models, thereby significantly simplifying the workflow and saving the cost. For instance, as shown in Figure 6, users can specify specific objects within images containing multiple elements through textual instructions (“the cat is the one in |image_2|”), and generate new images based on these instructions without needing preliminary operations such as image cropping. As shown in Figure 9, OmniGen can directly generate images based on the conditional information contained in the reference image, without loading another model to preprocess the reference image.
@@ 5.2 Reasoning Ability
We have explored the reasoning capabilities of the model and presented the results in Figure 11. As shown in the left half of Figure 11, when given an instruction without explicitly specifying the object, such as “Where can I wash my hands? Please help me find the right place in |image_1|”, the model can recognize image contents and infer that a sink is needed. Consequently, the model identifies and indicates the area of the sink in the image. This functionality creates potential applications in the field of embodied intelligence, assisting intelligent agents in comprehending multi-modal instructions, locating necessary objects and planning subsequent actions. Moreover, the right half of Figure 11 demonstrates that after inferring the target object, the model can also perform editing operations on it. If no object matches, the model will refrain from editing any unrelated objects.
Figure 11: Reasoning Ability of OmniGen
@@ 5.3 Chain of Thought
Figure 12: Step-by-Step Image Generation Illustration Simulating the Human Drawing Process
@@
6 Limitations and Discussions
Figure 13: Failure cases of OmniGen

### v2

@@ 3.2.3 Computer Vision Tasks
We introduce classic computer vision tasks to expand the boundaries of the model’s generative capabilities. For low-level vision tasks (low-light image enhancement [67], deraining [73], deblurring [44], inpainting [52], outpainting [52] and colorization [57]), where the annotation itself is an image, we only add text instructions, which are randomly sampled from instructions generated by GPT-4o. For high-level vision tasks, we choose to represent all annotations as images. We use images from LAION [57] as source images and annotations from [52] as targets to construct image pairs (such as source image and its human pose mapping). The annotations include human pose, depth mapping, canny, Hed edge, and segmentation. Additionally, we also use several datasets for referring image segmentation, including RefCOCO [26], ADE20k [78], and ReasonSeg [29]. As shown in Figure 3-(d), the input is the source image and a natural language expression, the output is an image with the corresponding object highlighted in blue.
@@ 3.3 Few-shot to Image
We build a few-shot to image dataset to stimulate the in-context learning ability of the model. Specifically, for each task described in the preceding sections, we randomly selected a few examples and combined the original input with these examples to form new inputs. The specific data format is shown in Figure 3-(e). Due to limitations in training resources, we opted to use only one example to improve training efficiency.
@@ 4 Experimental Results
The default inference step is set to 50. We use the classifier-free guidance for both text condition and image condition following  [3]. The default guidance scale is set to 2.5, and the image guidance scale is set to 1.6. For image editing task, we increase the image guidance scale to 2.
@@ 4.1.1 Qualitative Results
Figure 4: The results of different image generation tasks.
Figure 5: Subject-driven generation. When the reference image contains multiple objects, OmniGen can automatically identify the required objects based on textual instructions without the need for additional preprocessing steps like image cropping or face recognition.
Figure 4 summarizes the results of different image generation tasks, including text-to-image, image editing, and visual conditional. These results demonstrate that OmniGen can handle various downstream tasks based on multi-modal instructions.
Figure 5 presents the outcomes of the subject-driven generation task. OmniGen can extract the required objects from the given reference images and generate new images accordingly. Furthermore, when the reference image contains multiple objects, the model can directly select the needed objects based on textual instructions (e.g., the one wearing the dark red shirt in image) without requiring additional preprocessing steps. In contrast, existing models [64, 22] require manually cropping the target person from a multi-person image, detecting faces with a face detector, encoding faces with a face encoder, and inputting them into the diffusion model through a cross-attention module. Without these complex processes, OmniGen can complete the task in a single end-to-end step.
More qualitative results are provided in supplementary materials.
@@ 4.1.2 Text to Image
Following [12], we evaluate text-to-image generation capability on the GenEval [20] benchmark. We compare the performance of OmniGen with the reported results of other popular image generation models, as summarized in Table 2. Surprisingly, OmniGen achieves similar performance compared to the current diffusion models, such as SD3, which underscores the effectiveness of our framework. Notably, OmniGen has only 3.8 billion parameters, whereas the SD3 model has a total of 12.7 billion parameters. The architecture of OmniGen is significantly simplified, eliminating the cost of an additional encoder, thereby greatly enhancing the efficiency of parameter utilization.
@@
Model Params Overall object object Counting Colors Position binding
@@
IF-XL 5.5B+4.7B 0.61 0.97 0.74 0.66 0.81 0.13 0.35
SD3 8.0B+4.7B 0.68 0.98 0.84 0.66 0.74 0.40 0.43
OmniGen 3.8B 0.70 0.99 0.86 0.64 0.85 0.31 0.55
Table 2: Results of GenEval Benchmark. means the parameter of the frozen text encoder. Our model achieves comparable performance with a relatively small parameter scale.
4.1.3 Multi-modal to Image
@@
I-Pix2Pix [3] 0.219 0.834
MagicBrush [74] 0.222 0.838
PnP [63] 0.089 0.521
Null-Text [42] 0.236 0.761
EMU-Edit [58] 0.231 0.859
OmniGen 0.231 0.829
@@
Tex-Inv [13] 0.255 0.780
DreamBooth [56] 0.305 0.803
ReImagen [7] 0.270 0.740
SuTI [9] 0.304 0.819
Kosmos-G [47] 0.287 0.847
OmniGen 0.315 0.801
Table 3: Left: Results on EMU-Edit test data. Right: Results on DreamBench. As a universal model, OmniGen demonstrates performance comparable to that of the best proprietary models.
@@
We evaluate the image editing on EMU-Edit [58] dataset and subject-driven generation capability on DreamBench [56]. We use CLIP-T to measure how well the model followed the instructions, while CLIP-I similarity scores measure the model’s ability to preserve elements from the source image. As shown in Table 3, OmniGen exhibits comparable performance to the current state-of-the-art models. Noted for subject-driven generation task, we select one image of the specific object as input instead of fine-tuning the model like DreamBooth.
In Table 4, we use the dataset and script from  [31] to evaluate the generation capability based on visual conditions. For each condition, the controllability is evaluated by measuring the similarity between the input conditions and the extracted conditions from generated images of diffusion models. We can find that OmniGen achieves competitive results for all conditions.
Figure 6: The results of OmniGen in various traditional vision tasks.
Figure 7: ControlNet involves multiple steps and plugins: firstly using the corresponding detector to extract information from the reference image, and then loading the appropriate ControlNet module to process the visual conditions. In contrast, OmniGen completes the entire task in a single step within a single model.
@@ 4.2 Computer Vision Tasks
We present several qualitative results of computer vision tasks in Figure 6. OmniGen can handle various low-level vision tasks such as deraining, deblurring, and inpainting. At the bottom of Figure 6, we can see that OmniGen is also able to handle high-level vision tasks, such as human pose recognition.
Note that OmniGen’s capabilities in CV tasks are not intended to surpass existing state-of-the-art models that developed for specific tasks over a long period. We hope it can transfer the knowledge gained from these traditional computer vision tasks to image generation tasks, thereby unlocking greater potential. Some examples are shown in Figure 7. The existing workflow for ControlNet involves using a detector to extract spatial condition information from the reference image, and then loading the corresponding control module to model the spatial condition information for image generation. Can we directly use OmniGen to generate new images based on a reference image in only one step? Surprisingly, even without having encountered such a task before, OmniGen handles it admirably without explicit extraction of additional conditional images. More specifically, we can directly input the reference image and text instruction (e.g., Follow the human pose of this image to generate new image.) to generate the target image in only one step without any explicit additional intermediate procedures.
@@ 5 Further Analysis
LLMs demonstrate remarkable generalization capabilities, and can boost performance through mechanisms such as in-context learning and chain of thought. We observe similar functionalities in OmniGen as well, and present our findings in this section.
@@ 5.1 Emgerging Capabilities
By standardizing all tasks into a unified format and training on X2I dataset, OmniGen can acquire universal knowledge and allow knowledge transfer across different scenarios and tasks, thus enabling the generation capabilities on unseen tasks and domains.
@@
Task Composition. In real-world applications, user requirements often involve combinations of tasks. As shown in Figure 8-(a), our model is capable of simultaneously processing multiple instructions, including those for different tasks as well as multiple instructions for the same task. These results highlight our model’s versatility and potential for widespread adoption in the wild.
End-to-end Workflow. Users typically need to load multiple models and perform multi-step processing to ultimately generate a satisfactory image, making the workflow very cumbersome and costly. OmniGen possesses both excellent multi-modal understanding and image generation capabilities, enabling it to complete a lot of tasks without relying on external models, thereby significantly simplifying the workflow and saving cost. As demonstrated in Figure 7, OmniGen can extract the relevant conditional information from the reference image and generate a new image based on the captured condition within one step. This process is implicit, with all processing completed internally within the model, only requiring the user to input a simple command.
In-context Learning for Unseen Domains. We use the data from FSS [32] which contains objects that have never been seen in previous datasets to evaluate the generalized ability. In the right of Figure 8-(b), we can see that OmniGen is not familiar with the concepts of “pencil sharpeners”, and it cannot identify it from images. However, when provided with an example, the model is capable of making accurate predictions. By providing an example, the model is capable of successfully completing this unseen segmentation task.
@@ 5.2 Reasoning Ability
We have explored the reasoning capabilities of the model and presented the results in Figure 9. As shown in the left half of Figure 9, when given an instruction without explicitly specifying the object, such as “Where can I wash my hands? Please help me find the right place in image_1”, the model can recognize image contents and infer that a sink is needed. Consequently, the model identifies and indicates the area of the sink in the image. This functionality creates potential applications in the field of embodied intelligence, assisting intelligent agents in comprehending multi-modal instructions, locating necessary objects and planning subsequent actions. Moreover, the right half of Figure 9 demonstrates that after inferring the target object, the model can also perform editing operations on it. If no object matches, the model will refrain from editing any unrelated objects(see the example in the bottom right corner of Figure 9).
Figure 9: Reasoning ability of OmniGen.
@@ 5.3 Chain of Thought
Figure 10: Simulate the process of human drawing via step-by-step generation.
@@
6 Ablation Study
In this section, we assess the necessity of certain modules in OmniGen
Attention Mask. Figure 11-Top shows the generated images without the modified attention in Section 2.1. As we can see, there is a lot of noise in the images. Compared to Figure 4, the generated images in Figure 11-Top have many distorted areas, confirming that the unidirectional attention mechanism in LLMs is not suitable for the rectified flow.
Weighted Loss. Figure 11-Bottom shows the results when the weighted loss from Section 2.2 is not used for training. For image editing tasks, since the edited area is usually small, the model tends to learn to copy the input image directly as the output. Comparing Figure 11-Bottom and Figure 4, it’s evident that using weighted loss can prevent the model from learning this shortcut.
Input Image Representation. The CLIP model is widely used in multi-modal understanding models [37]. We also explored the difference between using CLIP [^2] https://huggingface.co/openai/clip-vit-large-patch14-336 and using VAE to process input images. The results are shown in Table 5. There is not much difference between the two strategies. Using VAE has a slight advantage in image similarity, which might be related to the fact that the output images also require VAE processing. Another advantage of using VAE is that it avoids additional modules, maintaining the simplicity of the model. Directly inputting the original image is another simpler approach that we leave for future exploration.
Figure 11: Top: results of causal attention; Bottom: results without loss weights.
Input Image EMU-Edit DreamBench
Representation CLIP-T CLIP-I CLIP-T CLIP-I
CLIP 0.232 0.820 0.312 0.789
VAE 0.231 0.829 0.315 0.801
Table 5: Comparison of different methods to obtain visual embedding for input images.
7 Related Work
Generative Foundation Models. The GPT series [53, 46] have demonstrated that language models can learn numerous tasks via training on a large-scale dataset. Beyond language, multi-modal large language models [37, 11] have been proposed to integrate vision and language capabilities. However, they lack the capability to generate images. Some works propose integrating LLMs with diffusion models to equip LLMs with image generation capability [60, 18, 69, 61]. Others use discrete tokens to support both image and text generation simultaneously [62, 41, 71]. They focus on multi-modal generation with limited image-generation capability. Concurrent works such as TransFusion [79] and Show-O [71] unify diffusion and autoregressive methods into a single model, generating text autoregressively and images through diffusion. Nonetheless, like existing diffusion models, these works focus on a limited range of image generation tasks, primarily text-to-image generation, and cannot cover more complex and various visual generation tasks. The construction of a universal foundation model for image generation remains unclear and has not been fully explored.
Diffusion Model. Recent advancements in diffusion models have been remarkable, with notable contributions from the Stable Diffusion (SD) series [55, 51, 12], DALL-E [54], and Imagen [25]. These models are predominantly designed for text-to-image generation tasks. Many efforts have been made to extend the capabilities of diffusion models, such as ControlNet [75], T2I-Adapter [43], StyleShot [15]. InstructPix2Pix [3], InstructSD [48] and EMU-edit [58] explore performing general image editing tasks through instructions. However, these methods are task-specific, extending the capabilities of SD by modifying the model architecture. In contrast, OmniGen is a model that natively supports various image-generation tasks, and no longer requires any preprocessing steps or assistance from other models. There is some work exploring the unification of computer vision (CV) tasks [2, 65, 14, 19]. However, these efforts primarily focus on classic vision tasks instead of general image generation tasks, and often underperform compared to those specifically designed and trained for corresponding tasks. In our work, the introduction of CV tasks does not aim to obtain a SOTA performance on CV tasks, which is designed to enable the model to learn general knowledge.
8 Limitations and Conclusion
In this work, we introduce OmniGen, the first unified image generation model. It is designed to be simple, flexible, and easy to use. We also construct the first large-scale unified image generation dataset, X2I, to activate the general capabilities of the OmniGen model. OmniGen has demonstrated excellent image-generation abilities across various tasks.
However, the current model still has some issues. The text rendering capability is limited, and long text can not be accurately generated. The output images may contain undesired details (e.g. abnormal hand). Besides, previously unseen image types (e.g., surface normal map) can hardly be processed as expected. More failure cases are provided in supplementary materials.
We believe that the future paradigm of image generation should be simple and flexible, allowing for the direct generation of various images through any multimodal instructions without complex workflows. OmniGen represents an important step towards a foundational model for universal image generation. We will open-source the relevant resources, hoping to provide insights for the future of image generation.
Supplementary Material
9 More Qualitative Results
In this section, we provide more qualitative results of OmniGen. Overall, OmniGen achieves competitive or even superior results compared to task-specific open-source models. It’s important to note that OmniGen is a general-purpose model and is not specifically trained for any particular task. Additionally, OmniGen has capabilities that many other models lack, allowing it to follow multimodal instructions to complete complex tasks. To the best of our knowledge, OmniGen is the first model capable of such flexible control in image generation.
9.1 Text to Image
Figure 12 shows more examples of text-to-image task. OmniGen can generate images with arbitrary aspect ratios.
9.2 Image Editing
In Figure 13, we provide qualitative results of OmniGen compared to some other open-source models. As shown, OmniGen achieves better results than existing open-source models, accurately modifying specified targets without altering other content. Emu-edit is not open-source, so further evaluation was not possible. Additionally, OmniGen demonstrates strong instruction comprehension capabilities, handling complex tasks such as executing two editing instructions in one step, as seen in Figure 13-(e), and reasoning editing tasks in the (f). In Figure 13-(f), we don’t not explicitly specify the target to be modified; the model infers the target (balloon) from the description and makes precise modifications. OmniGen inputs images and text into a unified transformer, allowing for extensive interaction between different modalities, resulting in powerful multimodal instruction following capabilities, which other frameworks like InstructPix2Pix cannot achieve.
9.3 Subject-driven
Figure 14 shows more examples of subject-driven generation on DreamBench. Similar to Kosmos-G, we use an image of an object as a reference and generate a new image based on the text description, without fine-tuning on the specific object. As illustrated, OmniGen achieves better subject fidelity and text fidelity, meaning it preserves the object from the reference image while following new textual instructions. Kosmos-G struggles to maintain the object from the reference image. Notably, Kosmos-G does not always adhere to the text instructions; as shown in Figure 14-(d), it fails to depict the described content. Instead, Kosmos-G nearly replicates the reference image, which is why it has a higher CLIP-I similarity (see Table 3 in the main paper). Additionally, Kosmos-G uses SD1.5 as the image model, resulting in lower image quality compared to OmniGen.
9.4 Identity-Preserving Generation
Figure 15 shows the results of Identity-Preserving generation task. OmniGen achieves better quality than instandID and comparable results to the latest model PULID-FLUX  [^3] https://huggingface.co/spaces/yanze/PuLID-FLUX. It should be noted that as a general model, OmniGen’s facial similarity is not yet on par with the latest specialized model, PULID-FLUX, which is an area for future improvement. Unlike these models, which require a face detector to detect faces and a face encoder to extract facial features, OmniGen does not need task preprocessing steps or additional plugins. It automatically completes tasks with input images and instructions. Moreover, OmniGen does not only focus on faces; it can capture other useful reference information in the input image, such as clothing. For example, in Figure 15-(a) and -(b), OmniGen tends to preserve the clothing of the person in the input image. Of course, when clothing is specified in the instructions, OmniGen can also follow the instructions to use new clothing (see Figure 15-(c) and -(d)). In contrast, other models can only use the extracted face and cannot utilize the original clothing from the input image.
Figure 16 showcases some unique features of OmniGen. Existing models cannot handle input images with multiple objects or multiple input images. OmniGen can extract desired individuals from a group photo based on instructions, and it can also process multiple images simultaneously, as shown in the Figure 16-(a). Besides, OmniGen can handle more than just people; as demonstrated in Figure 16-(b), it can identify specific clothing items in an image and use them to generate new images.
9.5 Failure cases